# The Separability Problem in Quantum Mechanics: Insights from Research on Axiomatics and Human Language

### v1

@@ 1 Introduction
Another reason why Einstein’s attention was focused on separability from the early 20th century onwards was his contributions in the formulation of the new (Bose-Einstein) quantum statistics and the fact it implied a lack of statistical independence of the microstates of identical quantum entities (Howard,, 1990). This lack of statistical independence, a consequence of quantum correlations, led Einstein to conjecture the existence of a mysterious force at the origin of the phenomenon, whose nature, however, remained to be clarified and remains an open question to this day. In Aerts et al., 2023b ; Aerts et al., 2023a ; Aerts et al., (2024), we proposed to understand this mysterious force, which brings together identical bosons, as the physical counterpart of the “meaning force” active in a text that tells a story, and the associated mechanism of contextual updating carried by meaning. This powerful analogy, between ‘quantum behavior’ and ‘cognitive behavior’, will be further explored in the final part of this article, to shed light on the question of the existence, or non-existence, of separate physical systems and the necessity of having theories that, from a structural point of view, can accomodate to describe such situations.
@@
Coming back to the EPR paper, Einstein was not happy with Podolsky formulation precisely because it failed to make explicit this central assumption of separability and, as Howard rightly points out (Howard,, 1985), this left the door open to the kind of criticism that Bohr made, for example pwhen in his response to the EPR article he wrote the following cryptic sentence (Bohr,, 1935): “[…] there is essentially the question of an influence on the very conditions which define the possible types of predictions regarding the future behavior of the system. Since these conditions constitute an inherent element of the description of any phenomenon to which the term ‘physical reality’ can be properly attached, we see that the argumentation of the mentioned authors does not justify their conclusion that quantum-mechanical description is essentially incomplete.” What Bohr refers to here is the fact that the system considered, with the associated experimental setting, is not to be considered a composite system formed by two separate sub-systems. So, again, the central element of the dispute between Einstein and Bohr is precisely separability.
@@
The development of this new mathematics required the pioneering work of Birkhoff and von Neumann (Birkhoff & von Neumann,, 1936), which gave rise to a new field of inquiry focused on being able to obtain an axiomatic foundation of quantum mechanics, possibly realistic and operational. This work found its greatest development in the 1970s and 1980s with the work of Mackey, Jauch, Piron, Foulis, Randall and Ludwig (Mackey,, 1963; Jauch,, 1968; Piron,, 1976; Foulis & Randall,, 1978; Ludwig,, 1983). One of the authors, a student of Constantin Piron, has been very actively involved in this quantum axiomatics research over the past century, so much so that the Geneva school of quantum mechanics has become known – at least among insiders – as the Geneva-Brussels school of quantum mechanics (Aerts,, 1982, 1986; Aerts, 1999b, ; Aerts & Durt,, 1994; Aerts et al.,, 1997; Aerts, Coecke & Smets,, 1999).
@@ 2 Separate systems
The first four of these seven axioms are very general and are connected with basic principles that one can expect to be satisfied for any type of system. On the other hand, the two “failing axioms”, those that are not satisfied by separate quantum systems are among the three axioms that had to be explicitly introduced to have the representation of the state space of quantum entities to become a standard Hilbert space. Different from the other four, they are rather technical mathematical axioms with no straightforward connection to physical intuition or known physic’s principles. That the two failing axioms are of a technical mathematical nature while the non-failing axioms are very general, e.g., similar to the axioms for set theory, is an important aspect in relation to the following question: Is Hilbert space too specific a mathematical structure to describe all the reality associated with quantum systems, or, simply, are there no separate quantum entities? We will ponder on this question in the last part of this article.
@@ 3 State, properties and the first axiom: completeness
Note also that, although we have described the completeness of the partially ordered set of properties, for the ‘stronger than’ partial order relation, as the first axiom, in fact, as we have seen, in the Geneva-Brussels formalism it is a theorem. The reason why it is generally mentioned as an axiom is because, in addition to the Geneva-Brussels formalism, these notions are also of interest to the broader field of quantum logic, for which, in the absence of an operational approach, the completeness of the lattice of properties is necessarily to be presented as an axiom.
@@ 5 Classical and quantum disjunctions
Due to the presence of entanglement, the conjunction measurement consisting of two yes-no measurements, $a$ and $b$, jointly performed by Alice and Bob, will produce anti-correlations, if both chose the same spatial direction, i.e., the experiment will give with certainty one of following two outcomes: (yes, no) or (no, yes), and both are of course equally possible. Indeed, if Alice’s spin is observed to be ‘up’, along a given direction, Bob’s spin is observed to be ‘down’, along that same direction, and vice versa, with the ‘up’ and ‘down’ outcomes playing here the role of the ‘yes’ and ‘no’ answers. Now, the logical way to define a disjunction property using a conjunction measurement is to consider that the property is actual (answer ‘yes’) if Alice’s test, or Bob’s test, or both, give the answer ‘yes’, and potential otherwise. In other words, the three outcomes (yes, yes), (yes, no) and (no, yes) are associated to the answer ‘yes’, for the yes-no measurement testing the disjunction property $a\vee b$, and the outcome (no, no) is associated with the answer ‘no’.
@@ 9 Covering law, weak modularity and plane transitivity
Regarding the sixth axiom, weak modularity, what we can say is that, on the level of the Hilbert space, it is the equivalent of the Cauchy completeness of that Hilbert space, if the formalism is represented within a real, complex or quaternionic vector space structure. In other words, it is the requirement that all converging sequences, in the Cauchy sense, converge within the state space and that there are no “missing elements.” In technical terms, the axiom requires that if $a$ and $b$ are two properties such that $a<b$, then $(a\vee b^{\prime})\wedge b=a$.
@@ 10 The violated axioms
Purely operational axiomatic work probably lies ahead as well. Which structure can be identified as operationally inferable and therefore requires less structure to be added from an axiomatic standpoint is not a fully resolved question at present, as it is the case with which classical structures can be identified as substructures, operationally or axiomatically. An example is how we managed to operationally introduce the notion of ‘topological’ (considering topological as the classical for closure) as an alternative to ‘classical’ (in the sense of deterministic) and characterize the corresponding substructures in citeaertsdhooghesioen2011, with the intention of making a pass towards the generally continuous (hence topological) formalisms of classical mechanics.
@@ 11 Separability in human language
If this correspondence between quantum physical entities and conceptual entities is not merely accidental, but an expression of the nature of our physical reality, at a fundamental level, it is also possible that the behaviour of quantum physical entities can also be explained by assuming that their nature is not particle-like, or wave-like, or something intermediary between a particle and a wave, but conceptual, in the sense that our physical world would ultimately, at a fundamental level, be governed by dynamics which depend on the meaning that the different physical entities would carry. According to the conceptuality interpretation elaborated by one of us, the duality between bosonic and fermionic entities would then be, for example, a reflection of the duality between conceptual entities, which carry meaning, and cognitive entities, which are sensitive to meaning (Aerts, 2009b, ; Aerts et al.,, 2020).
We cannot, of course, go into the merits of this perspective on reality, where human language and the corresponding cognitive processes are seen as just one of many cognitive-conceptual structures that would have emerged in the course of evolution, from the very first moments of existence of our universe (Aerts & Sozzo,, 2015; Aerts & Sassoli de Bianchi,, 2018, 2022). We do not know whether such a perspective will find some compelling experimental confirmation in the future. At present, it remains a speculative hypothesis that has the advantage of offering numerous powerful explanations for the otherwise incomprehensible behavior of microscopic quantum systems. Here, we will use it to help us reason about the separability issue, aware that it may turn out to be much more than just a suggestive analogy, in the future.

### v2

@@ 1 Introduction
Another reason why Einstein’s attention was focused on separability from the early 20th century onwards was his contributions in the formulation of the new (Bose-Einstein) quantum statistics [^3] The Bose–Einstein statistics is a procedure for counting the states of systems composed of indistinguishable entities with integer spin. and the fact it implied a lack of statistical independence of the microstates of identical quantum entities (Howard,, 1990). This lack of statistical independence, a consequence of quantum correlations, [^4] Quantum mechanics predicts that joint measurements on composite systems in entangled states reveal correlations between observed values. The pattern exhibited by these correlations, however, seems to conflict with classical intuitions, particularly the idea that there may be a common cause in the past at the origin of the correlations. These seem instead to indicate that a composite system in an entangled state behaves more as an indivisible whole than as two separate subsystems, and that the observed correlations are created and not discovered by the act of observation. led Einstein to conjecture the existence of a mysterious force at the origin of the phenomenon, whose nature, however, remained to be clarified and remains an open question to this day. In Aerts et al., 2023b ; Aerts et al., 2023a ; Aerts et al., (2024), we proposed to understand this mysterious force, which brings together identical bosons, [^5] In physics, bosons are entities obeying the Bose–Einstein statistics. Some elementary bosons act as force carriers, like photons for the electromagnetic force and gluons for the strong nuclear force. as the physical counterpart of the “meaning force” active in a text that tells a story, and the associated mechanism of contextual updating carried by meaning. This powerful analogy, between ‘quantum behavior’ and ‘cognitive behavior’, will be further explored in the final part of this article, to shed light on the question of the existence, or non-existence, of separate physical systems and the necessity of having theories that, from a structural point of view, can accomodate to describe such situations.
@@
Coming back to the EPR paper, Einstein was not happy with Podolsky’s formulation precisely because it failed to make explicit this central assumption of separability and, as Howard rightly points out (Howard,, 1985), this left the door open to the kind of criticism that Bohr made, for example pwhen in his response to the EPR article he wrote the following cryptic sentence (Bohr,, 1935): “[…] there is essentially the question of an influence on the very conditions which define the possible types of predictions regarding the future behavior of the system. Since these conditions constitute an inherent element of the description of any phenomenon to which the term ‘physical reality’ can be properly attached, we see that the argumentation of the mentioned authors does not justify their conclusion that quantum-mechanical description is essentially incomplete.” What Bohr refers to here is the fact that the system considered, with the associated experimental setting, is not to be considered a composite system formed by two separate sub-systems. So, again, the central element of the dispute between Einstein and Bohr is precisely separability.
@@
The development of this new mathematics required the pioneering work of Birkhoff and von Neumann (Birkhoff & von Neumann,, 1936), which gave rise to a new field of inquiry focused on being able to obtain an axiomatic foundation of quantum mechanics, possibly realistic and operational. This work found its greatest development in the 1970s and 1980s with the work of Mackey, Jauch, Piron, Foulis, Randall and Ludwig (Mackey,, 1963; Jauch,, 1968; Piron,, 1976; Foulis & Randall,, 1978; Ludwig,, 1983). One of the authors, a student of Constantin Piron, has been very actively involved in this quantum axiomatics research over the past century, so much so that the Geneva school of quantum mechanics has become known – at least among insiders – as the Geneva-Brussels school of quantum mechanics (Aerts,, 1982, 1986; Aerts, 1999b, ; Aerts & Durt,, 1994; Aerts et al.,, 1997; Aerts, Coecke & Smets,, 1999). [^6] One should not think of the Geneva-Brussels School as a physical institution where formal education or research takes place, but rather as an intellectual tradition in which scholars pursue a common approach in exploring the foundations of physical theories and, more recently, of human cognition. The name is primarily due to the pioneering contributions of Ernst Stückelberg, Jospeh Maria Jauch and Constantin Piron, and their collaborators at the University of Geneva, and the subsequent developments achieved by Diederik Aerts and his collaborators at the Vrije Universiteit Brussel.
@@ 2 Separate systems
The first four of these seven axioms are very general and are connected with basic principles that one can expect to be satisfied for any type of system. On the other hand, the two “failing axioms”, those that are not satisfied by separate quantum systems are among the three axioms that had to be explicitly introduced to have the representation of the state space of quantum entities to become a standard Hilbert space. Different from the other four, they are rather technical mathematical axioms with no straightforward connection to physical intuition or known physical principles. That the two failing axioms are of a technical mathematical nature while the non-failing axioms are very general, e.g., similar to the axioms for set theory, is an important aspect in relation to the following question: Is Hilbert space too specific a mathematical structure to describe all the reality associated with quantum systems, or, simply, are there no separate quantum entities? We will ponder on this question in the last part of this article.
@@ 3 State, properties and the first axiom: completeness
Note also that, although we have described the completeness of the partially ordered set of properties, for the ‘stronger than’ partial order relation, as the first axiom, in fact, as we have seen, in the Geneva-Brussels formalism it is a theorem. The reason why it is generally mentioned as an axiom is because, in addition to the Geneva-Brussels formalism, these notions are also of interest to the broader field of quantum logic, [^10] Quantum logic is a formal logical system that considers propositions as subspaces of a Hilbert space. This results in a specific calculus of propositions different from classical logic, in that some well-known principles, like the law of distributivity, no longer apply. Quantum logic can nevertheless be considered a genuine logic if we take logic as a tool for structuring reasoning within a specific domain. From this point of view, it provides a coherent system for reasoning about quantum entities, just as classical logic provides a coherent system for reasoning about macroscopic and deterministic entities. for which, in the absence of an operational approach, the completeness of the lattice of properties is necessarily to be presented as an axiom.
@@ 5 Classical and quantum disjunctions
Due to the presence of, entanglement, [^12] In quantum mechanics, entanglement is a situation in which a composite system is in a state that contains no information about the (pure) states of its subsystems. This is mathematically expressed as a non-factorizability of the state of the composite system, being a superposition involving both subsystems, at the origin of quantum correlations. the conjunction measurement consisting of two yes-no measurements, $a$ and $b$, jointly performed by Alice and Bob, will produce anti-correlations, if both chose the same spatial direction, i.e., the experiment will give with certainty one of following two outcomes: (yes, no) or (no, yes), and both are of course equally possible. Indeed, if Alice’s spin is observed to be ‘up’, along a given direction, Bob’s spin is observed to be ‘down’, along that same direction, and vice versa, with the ‘up’ and ‘down’ outcomes playing here the role of the ‘yes’ and ‘no’ answers. Now, the logical way to define a disjunction property using a conjunction measurement is to consider that the property is actual (answer ‘yes’) if Alice’s test, or Bob’s test, or both, give the answer ‘yes’, and potential otherwise. In other words, the three outcomes (yes, yes), (yes, no) and (no, yes) are associated to the answer ‘yes’, for the yes-no measurement testing the disjunction property $a\vee b$, and the outcome (no, no) is associated with the answer ‘no’.
@@ 9 Covering law, weak modularity and plane transitivity
Regarding the sixth axiom, weak modularity, what we can say is that, on the level of the Hilbert space, it is the equivalent of the Cauchy completeness of that Hilbert space, if the formalism is represented within a real, complex or quaternionic vector space structure. In other words, it is the requirement that all converging sequences, in the Cauchy sense, converge within the state space and that there are no missing elements. [^18] More precisely, Cauchy completeness ensures that “gaps” do not exist in the space in question, in the sense that Cauchy sequences always converge to elements in it. For example, the set of real numbers with the usual distance metric is Cauchy complete, but the set of rational numbers is not, as there are Cauchy sequences of rational numbers that converge to irrational numbers. In technical terms, the axiom requires that if $a$ and $b$ are two properties such that $a<b$, then $(a\vee b^{\prime})\wedge b=a$.
@@ 10 The violated axioms
Purely operational axiomatic work probably lies ahead as well. Which structure can be identified as operationally inferable and therefore requires less structure to be added from an axiomatic standpoint is not a fully resolved question at present, as it is the case with which classical structures can be identified as substructures, operationally or axiomatically. An example is how we managed to operationally introduce the notion of ‘topological’ (considering topological as the classical for closure) as an alternative to ‘classical’ (in the sense of deterministic) and characterize the corresponding substructures in citetaertsdhooghesioen2011, with the intention of making a pass towards the generally continuous (hence topological) formalisms of classical mechanics.
@@ 11 Separability in human language
If this correspondence between quantum physical entities and conceptual entities is not merely accidental, but an expression of the nature of our physical reality, at a fundamental level, it is also possible that the behaviour of quantum physical entities can also be explained by assuming that their nature is not particle-like, or wave-like, or something intermediary between a particle and a wave, but conceptual, in the sense that our physical world would ultimately, at a fundamental level, be governed by dynamics which depend on the meaning that the different physical entities would carry. According to the conceptuality interpretation elaborated by one of us, the duality between bosonic and fermionic entities [^20] Unlike bosons, which obey the Bose–Einstein statistic and are integer spin entities, fermions obey the Fermi–Dirac statistic and are fractional spin entities. In particular, they obey the Pauli exclusion principle, which gives stability to matter by preventing identical fermions from occupying the same quantum state. would then be, for example, a reflection of the duality between conceptual entities, which carry meaning, and cognitive entities, which are sensitive to meaning Aerts, 2009b ; Aerts et al., (2020).
We cannot, of course, go into the merits of this perspective on reality, where human language and the corresponding cognitive processes are seen as just one of many cognitive-conceptual structures that would have emerged in the course of evolution, from the very first moments of existence of our universe (Aerts & Sozzo,, 2015; Aerts & Sassoli de Bianchi,, 2018, 2022). [^21] The conceptuality interpretation did not arise from a fundamental philosophical principle, as is often the case with other interpretations, but from the analysis of concrete paradoxical quantum situations and the attempts to make sense of them. Fundamental to the interpretation is the insight that if bosons, the force-carrying quantum particles, are not objects but conceptual entities, then these paradoxical situations suddenly become intelligible. In this sense, the conceptuality interpretation is a bottom-up interpretation that proposes a radically new ontology for quantum entities. This makes it difficult to compare with the other existing quantum interpretations, none of which, in our view, provides a good explanation for all known quantum phenomena.
We do not know whether such a perspective will find some compelling experimental confirmation in the future. At present, it remains a speculative hypothesis that has the advantage of offering numerous powerful explanations for the otherwise incomprehensible behavior of microscopic quantum systems. Here, we will use it to help us reason about the separability issue, aware that it may turn out to be much more than just a suggestive analogy, in the future.